suppose you are canoeing along a river with a steady current. your average speed upstream is 2.5 mi/h. on the return trip you paddle with the current, and your average speed is 4 mi/h. find the average speed of the current and your average speed if you were paddling in still water

Answer by mananth(16946)

speed of canoe in still water = x mph
speed of current = y mph
Upstream speed = x-y =2.5
downstream speed = x+y = 4
add both equations
2x=6.5
x= 3.25
y= 3.25 -2.5 = 0.75
speed of canoe in still water = 3.25 mph
speed of current = 0.75 mph